Solve for x of the line through the two given points has the given slope.
(1ˏ−1)and(xˏx^2−2), m=4
I just don't understand how to find x.
I got the equation y=4x-5 with what was given but I don't know how that helps.

Answer by ewatrrr(24785)

 
Hi,
Line: y =4x-5 Good Work
4x-5 = x^2 - 2
x^2 - 4x + 3) =0
(x-1)(x-3) = 0, x = 1 and x = 3
Note: P(3,7)